The length of rectangle is 2 inches less than twice its width.if the area of the rectangle is 112 square inches.find the length of the shorter leg.
Answer by Jeetbhatt10th(11)

Let width be x units
therefore length=2x-2
Since area = length×breadth
Therefore 112=(2x-2)(x)
2x^2-2x=112
x^2-x=56
x^2-x-56=0
x^2-8x+7x-56=0
x (x-8)+7 (x-8)=0
(x+7)(x-8)=0
=>x=8,-7
But length cannot be negative,thus x=8
Width=8 inches
Length=2x-2=2×8-2=14 inches
Hence dimensions are 14 inches and 8 inches (Answer)☺